How would factor these trinomial factors 9z^2-64?
Answer by Theo(13342)

a^2 - b^2 = (a-b) * (a+b)

that's the general form.

your equation is:

9z^2 - 64 = (3z)^2 - 8^2

this is the same form if you let (3z) = a and 8 = b

you get 9z^2 - 64 = (3z)^2 - 8^2 = (3z - 8) * (3z + 8)

multiply those out and you get (3z)^2 - 3z*8 - 3z*8 - 8^2 which becomes (3z)^2 - 8^2 which becomes 9z^2 - 64.